Solve lnx= 3+ln(x-3)
Answer by stanbon(48535)

Solve lnx= 3+ln(x-3)
lnx-ln(x-3) = 3
ln[x/(x-3)] = 3
x/(x-3) =e^3
x/(x-3) = 20.0855...
x = (20.0855..)x - 60.2566..
19.0855x = 60.2566...
x = 3.15718...
Cheers,
Stan H.